How do I do the following 2 problems related to parallels and perpendiculars?
"Write an equation of the line through the point (a) parallel to the given line and (b) perpendicular to the given line". I must show my work. I am a 53 year old non-traditional student, and this algebra is really difficult for me to understand. Thank you so much for your time, I really appreciate it! Have a blessed day...Kathye Watson
1. Point = (3, -2) Line = 5x - 4y = 8
a)parallel = ? b)perpendicular = ?

Answer by richwmiller(17219)

parallel has the same slope m
perpendicular slope m/-1
1-slope = -a/b -5/-4=5/4
parallel 5/4
3,-2
y=5/4x+b
-2=5/4*3+b
-8/4-15/4=b
-23/4=b
y=5x/4-23/4
or
4y=5x-23
perp -4/5
y=-4/5+b
-2=-4/5*3+b
-10/5+12/5=b
2/5=b
y=-4/5x+2/5
or
5y=-4x+2
or
4x+5y=2
Now you get to try it on
2. Point = (-8, 3) Line = 2x + 3y = 5
a)parallel = ? b)perpendicular = ?
first get the slope -a/b a=2 b=3
NB this b is not the same as the b of y=mx+b
this b comes from ax+by+c=0